OpenStudy (anonymous):

-2y+3(y-1) when Y= -1 please solve

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Just plug in -1 for Y and solve -2(-1) + 3(-1-1)=... 2 + 3(-2)=... 2 - 6=-4

OpenStudy (anonymous):

thank you!